OpenStudy (ledah):

What is lateral area of a rectangular prism whose base is 6cm x 8cm and whose height is 5cm?

OpenStudy (anonymous):

la=2h(w+l) =2(5)(6+8) =10*14 =140

Parth (parthkohli):

Lateral area of a rectangular prism is \( \color{Black}{ 2h(w + l) }\)

Parth (parthkohli):

Can you substitute?

Parth (parthkohli):

First, do something like \(2(5)(6 + 8)\)

Parth (parthkohli):

Do you follow?

OpenStudy (ledah):

Yeah I follow

Parth (parthkohli):

Can you tell me what 2(5) is?

OpenStudy (ledah):

That would be 10

Parth (parthkohli):

Sorry, I was busy. Yes, 10. Now what is 8 + 6?

OpenStudy (ledah):

14

Parth (parthkohli):

What is 10 x 14?

OpenStudy (ledah):

140

Parth (parthkohli):

That's it.